### v1

# 3D Scene Prompting for Scene-Consistent Camera-Controllable Video Generation

Abstract
We present 3DScenePrompt, a framework that generates the next video chunk from arbitrary-length input while enabling precise camera control and preserving scene consistency. Unlike methods conditioned on a single image or a short clip, we employ dual spatio-temporal conditioning that reformulates context-view referencing across the input video. Our approach conditions on both temporally adjacent frames for motion continuity and spatially adjacent content for scene consistency. However, when generating beyond temporal boundaries, directly using spatially adjacent frames would incorrectly preserve dynamic elements from the past. We address this by introducing a 3D scene memory that represents exclusively the static geometry extracted from the entire input video. To construct this memory, we leverage dynamic SLAM with our newly introduced dynamic masking strategy that explicitly separates static scene geometry from moving elements. The static scene representation can then be projected to any target viewpoint, providing geometrically-consistent warped views that serve as strong 3D spatial prompts while allowing dynamic regions to evolve naturally from temporal context. This enables our model to maintain long-range spatial coherence and precise camera control without sacrificing computational efficiency or motion realism. Extensive experiments demonstrate that our framework significantly outperforms existing methods in scene consistency, camera controllability, and generation quality.
Figure 1: Teaser. Our framework generates the next video chunk that follows a user-specified camera trajectory while maintaining scene consistency. Our dual spatio-temporal conditioning combines the last few frames for temporal continuity and the rendered point cloud for spatial consistency.
1 Introduction
Camera-controllable video generation (He et al., 2024; Wang et al., 2024b; Jin et al., 2025) aims to synthesize videos following user-specified camera trajectories while maintaining visual coherence and temporal consistency. Recent advances have progressed from generating entirely new videos with controllable viewpoints (Bahmani et al., 2025) to enabling users to extend a single image or short video clips along desired camera paths (He et al., 2024; Agarwal et al., 2025). Yet these methods share a fundamental limitation: they can only process extremely short conditioning sequences, typically just a few frames, which constrains their ability to understand longer videos and hence fails to preserve the rich scene context present in those longer videos. What if we could provide a model with arbitrary-length video sequences and generate continuations that not only follow precise camera controls but also maintain scene consistency with the entire input? Such task, which we refer to as scene-consistent camera-controllable video generation, has immediate applications in film production (Zhang et al., 2025), virtual reality (He et al., 2025), and synthetic data generation (Knapp & Bohacek, 2025).
Scene-consistent, camera-controllable video generation presents three coupled challenges: first, static elements must remain stable across time, while dynamic elements e.g., moving objects or people, should evolve naturally from their most recent states rather than replaying motions from the distant past.Second, effective camera control requires understanding the scene’s 3D geometry: generated content must obey physical constraints, handle occlusions correctly, seamlessly compose dynamic elements with static structure, and plausibly extrapolate previously unobserved regions. Third, these capabilities must remain computationally practical: naïvely conditioning on all input frames scales poorly and becomes intractable as sequence length grows.
Rather than building a bespoke architecture, we leverage strong pretrained video generators, retaining their learned priors and training efficiency, by redesigning how prior content is referenced. Our key insight is to fundamentally rethink this referencing via a dual spatio-temporal conditioning mechanism. Current image-to-video (Yang et al., 2024) and video-to-future-video models  [^1] Throughout our paper, video-to-future-video models refer to models that are capable of generating the subsequent frames of the given input video (e.g., cosmos-predict2 (Agarwal et al., 2025). (Agarwal et al., 2025) achieve realistic generation by conditioning on temporally adjacent frames to maintain short-term consistency and motion continuity. However, adjacency in video is not purely temporal—it can also be spatial. When generating scene-consistent videos, the frames we synthesize may be spatially adjacent to frames from much earlier in the input sequence, particularly when the camera revisits similar viewpoints or explores nearby regions. This dual nature of adjacency suggests a new conditioning paradigm that leverages both temporal and spatial relationships.
Based on these motivations, we propose 3DScenePrompt, a novel video generation framework designed for scene-consistent camera-controllable video synthesis. It takes an arbitrary-length video as context and generates the future video that is consistent with the scene geometry of the context video. The key innovation lies in our dual spatio-temporal conditioning strategy: the model conditions on both temporally adjacent frames (for motion continuity) and spatially adjacent frames (for scene consistency). However, an important consideration for spatial conditioning for our task is that it must provide only the persistent static scene structure while excluding dynamic content, as directly conditioning on spatially adjacent frames from the past would incorrectly preserve dynamic elements. To enable this without temporal contradictions, we construct a 3D scene memory that exclusively represents the static geometry extracted from the entire input video.
To construct this 3D scene memory from dynamic videos, we leverage recent advances in dynamic SLAM frameworks (Zhang et al., 2022; 2024; Li et al., 2024) to estimate camera poses and 3D structure from the input video. To extract only the static regions from the estimated 3D structure, we introduce a dynamic masking strategy that explicitly separates static elements and moving objects. The static-only 3D representation can then be projected to target viewpoints, yielding geometrically-consistent warped views that serve as spatial prompts while allowing dynamic elements to evolve naturally from temporal context alone. Surprisingly, the integration of 3D scene memory provides an additional benefit: the geometrically-consistent warped views provide rich visual references that significantly reduce uncertainty in viewpoint manipulation, enabling precise camera control without any other explicit camera conditioning.
In summary, 3DScenePrompt enables both accurate camera control and long-range spatial consistency by treating the static scene representation as a persistent spatial prompt that guides generation across arbitrary timescales. Extensive experiments demonstrate that our framework significantly outperforms existing methods in maintaining scene consistency, achieving precise camera control, and generating high-quality videos from arbitrary-length inputs.
Figure 2: Comparison to existing architectures. (a) Camera-controllable methods condition on a single frame and camera trajectory (Wang et al., 2024b; He et al., 2024; Jin et al., 2025). (b) Video-to-future-video methods use the last few frames of the input video when generating the future video for temporal continuity (Song et al., 2025), but fail to maintain long-term spatial consistency when revisiting viewpoints unseen in the given few frames. (c) Our approach combines temporal conditioning (last few frames) with spatial conditioning (spatially adjacent frames) to achieve scene-consistent generation with precise camera control.
2 Related Work
Camera-controllable video generation.
Building upon the recent success of video diffusion models (Blattmann et al., 2023; Guo et al., 2023; Yang et al., 2024; Runway, 2024; Brooks et al., 2024), recent works (He et al., 2024; Wang et al., 2024b; Bahmani et al., 2024) have achieved camera-controllable video generation by introducing additional adapters into U-Net-based video diffusion models that accept camera trajectories. For instance, CameraCtrl and VD3D (Bahmani et al., 2024; He et al., 2024) incorporate spatiotemporal camera embeddings, such as Plücker coordinates, via ControlNet-like mechanisms (Zhang et al., 2023). While these methods enable precise trajectory following, they only condition on single starting images, lacking mechanisms to maintain consistency with extended video context. In contrast, our approach enables leveraging entire video sequences as spatial prompts through 3D memory construction, enabling scene-consistent generation that preserves the rich scene context within arbitrary-length inputs.
Geometry-grounded video generation.
Recent works (Ren et al., 2025; Yu et al., 2025; Seo et al., 2025) have integrated off-the-shelf geometry estimators into video generation pipelines to improve geometric accuracy. Gen3C (Ren et al., 2025), for instance, similarly adopts dynamic SLAM (Li et al., 2024; Han et al., 2025) to lift videos to 3D representations. However, these methods exclusively address dynamic novel view synthesis—generating new viewpoints within the same temporal window as the input. This constrained setting allows them to simply warp entire scenes without distinguishing static and dynamic elements. Our work fundamentally differs by generating content beyond temporal boundaries, requiring selective masking of dynamic regions during 3D construction—a critical challenge that emerges only when static geometry must persist while dynamics evolve naturally into the future.
Long-horizon scene-consistent generation.
Various approaches attempt scene-consistent long video generation through different strategies. ReCamMaster (Bai et al., 2025) and TrajectoryCrafter (Yu et al., 2025) interpolate frames or construct 3D representations but remain confined to the input’s spatiotemporal coverage, essentially performing dynamic novel view synthesis. StarGen (Zhai et al., 2025) scales to long trajectories but assumes static worlds, eliminating temporal dynamics entirely. DFoT (Song et al., 2025) most closely relates to our work, proposing guidance methods that condition on previous frames for scene consistency. However, DFoT also faces fundamental memory constraints when processing extended sequences, limiting its ability to maintain long-range spatial coherence. Our dual spatio-temporal strategy with SLAM-based spatial memory overcomes these limitations by selectively retrieving only the most relevant frames, both temporally and spatially, enabling computationally efficient processing of arbitrary-length videos while maintaining both motion continuity and scene consistency.
3 Methodology
3.1 Problem Formulation and Motivation
We address the task of scene-consistent camera-controllable video generation: given a dynamic video $\mathbf{V}_{\text{in}}\in\mathbb{R}^{L\times H\times W\times 3}$ of arbitrary length $L$ as context with height $H$ and width $W$, our goal is to generate $T$ subsequent frames $\mathbf{V}_{\text{out}}\in\mathbb{R}^{T\times H\times W\times 3}$ that follow a desired camera trajectory $\mathbf{C}=\{C_{t}\}_{t=1}^{T}$ while maintaining consistency with the scene captured in the context input:
$\mathbf{V}_{\text{out}}=\mathcal{F}(\mathbf{V}_{\text{in}},\mathcal{T},\mathbf{C}),$ (1)
where $C_{t}\in\mathbb{SE}(3)$ represents camera extrinsic matrices and $\mathcal{T}$ is a text prompt when a video generator $\mathcal{F}(\cdot)$ is based on pretrained text-to-video priors (Hong et al., 2023; Yang et al., 2024; Bahmani et al., 2025).
Comparison to existing solutions.
This task fundamentally differs from existing video generation paradigms. Existing camera-controllable generation methods (He et al., 2024; Wang et al., 2024b; Bahmani et al., 2024) synthesize videos following user-specified trajectories but only condition on a single image $\mathbf{I}_{\text{ref}}$ or plain text $\mathcal{T}$ (Fig. 2-(a)):
$\mathbf{V}_{\text{out}}=\mathcal{F}(\mathbf{I}_{\text{ref}},\mathcal{T},\mathbf{C}),\;\;\text{or}\;\;\mathbf{V}_{\text{out}}=\mathcal{F}(\mathcal{T},\mathbf{C}),$ (2)
which is insufficient for our task, where the entire underlying 3D scene of the context video should be considered. In contrast, video-to-future-video generation methods such as Cosmos-predict-2 (Agarwal et al., 2025) $\mathcal{G}(\cdot)$ employ temporal sliding windows to generate future frames (Fig. 2-(b)):
$\mathbf{V}_{\text{out}}=\mathcal{G}(\mathbf{V}_{\text{in}}[L-w:L],\mathcal{T})$ (3)
where $\mathbf{V}_{\text{in}}[L-w:L]$ for $w\ll L$ represents a small overlap window, typically consisting of the last few frames of $\mathbf{V}_{\text{in}}$. Although this design encourages temporal smoothness by providing the last few frames when generating the future video, it often fails to preserve long-term spatial consistency when the camera revisits regions not covered by the small window $w$.
Figure 3: Overview of 3DScenePrompt framework. To generate the next chunk video that remains spatially consistent with the input video, we design a dual spatio-temporal conditioning pipeline to extract the most relevant information from the input video. The last few frames are utilized to provide temporal conditioning, ensuring motion continuity between conditioning inputs and the generated frames. In parallel, for the spatial conditioning, we first select the most representative frames from the input sequence, lift their static regions into a 3D point cloud using the dynamic mask, and render it along a user-specified camera trajectory to preserve scene geometry.
3.2 Towards Scene-Consistent Camera-Controllable Video Generation
The key challenge of scene-consistent camera-controllable video generation lies in reconciling two competing requirements: maintaining consistency with potentially distant frames that share spatial proximity (when the camera returns to similar viewpoints), while evolving dynamic content naturally from the recent temporal context. Ideally, extending the conditioning window of video-to-future-video frameworks to condition on all frames $\mathbf{V}_{\text{in}}$ would ensure optimal global spatial consistency and dynamic smoothness. However, this quickly becomes impractical as the sequence grows, since standard self-attention incurs quadratic time/memory in the sequence length.
Dual spatio-temporal sliding window strategy.
Instead of increasing the temporal window size $w$ of the existing video-to-future-video generation methods, we introduce a dual sliding window strategy that conditions on frames selected along both temporal and spatial axes (Fig. 2-(c), and Fig. 3). Beyond the standard temporal window that captures recent motion dynamics, we add a spatial window that retrieves frames sharing similar 3D viewpoints, regardless of their temporal distance:
$\mathbf{V}_{\text{out}}=\mathcal{F}(\tilde{\mathbf{V}}_{\text{in}},\mathcal{T},\mathbf{C}),\;\;\text{where}\;\ \tilde{\mathbf{V}}_{\text{in}}=\{\text{Temporal}(w)\}\cup\{\text{Spatial}(T)\},$ (4)
where the model $\mathcal{F}$ generates a future sequence $\mathbf{V}_{\text{out}}$ conditioned on $\text{Temporal}(w)$, last $w$ frames of the input video $\mathbf{V}_{\text{in}}[L-w:L]$, and $\text{Spatial}(T)$, the $T$ retrieved frames from the entire input sequence based on viewpoint similarity to the target viewpoint $\mathbf{C}$. This dual conditioning enables the model to reference distant frames that observe the same spatial regions, maintaining scene consistency without processing all $L$ input frames.
While this dual conditioning is conceptually appealing, naïvely retrieving and providing spatially adjacent frames directly would be problematic for our task. Since we aim to generate content beyond the input’s temporal boundary, directly conditioning on much earlier frames would freeze outdated dynamics—for example, a walking person seen many frames ago should not reappear at the same location when synthesizing far-future frames. The spatial conditioning should therefore provide only the persistent scene structure while excluding dynamic content. Rather than retrieving individual frames, we introduce a 3D scene memory $\mathcal{M}$ that represents exclusively the static geometry extracted from all spatially relevant frames.
3.3 3D Scene Memory Construction
Our 3D scene memory must efficiently encode spatial relationships across all $L$ frames while extracting only persistent static geometry. To construct the 3D scene memory, we leverage dynamic SLAM frameworks (Li et al., 2024; Zhang et al., 2024) to estimate camera poses and reconstruct 3D structure:
$(\hat{\mathbf{C}},\mathbf{P})=\mathcal{D}_{\text{SLAM}}(V_{\text{in}}),$ (5)
where $\hat{\mathbf{C}}=\{\hat{C}_{i}\}_{i=1}^{L}$ are the estimated camera poses, $\mathbf{P}$ represents the aggregated 3D point cloud from the $L$ input frames, and $\mathcal{D}_{\text{SLAM}}(\cdot)$ represents the dynamic SLAM framework. This SLAM integration is effective in that it not only estimates the camera parameters of the input frames but also reconstructs the 3D structure of the scene, which can be further utilized to represent the 3D static geometry in a more efficient manner than other representations (Hong et al., 2024a; b).
While the camera poses $\hat{\mathbf{C}}$ enable efficient spatial retrieval by comparing viewpoint similarity with the target trajectory $\mathbf{C}$, the aggregated 3D point cloud $\mathbf{P}$ still contains both static and dynamic regions. Thus, we now explain our full pipeline on how to identify dynamic regions and only maintain the persistent static geometry of the input video.
Dynamic masking for static scene extraction.
Figure 4: Illustration of dynamic masking for static scene extraction. When aggregating 3D points across frames, moving objects create ghosting artifacts if not properly masked. (a) Without masking, dynamic elements (horses and riders) appear frozen at multiple positions, severely degrading the warped views. (b) With our dynamic masking pipeline, these elements are identified and excluded, resulting in clean static-only point clouds that can be reliably warped to new viewpoints.
Naïvely aggregating points across frames creates ghosting artifacts where moving objects appear frozen at multiple positions, as shown in Fig. 4-(a). We address this through a comprehensive three-stage masking pipeline that identifies and excludes all dynamic content as depicted in Fig. 5.
We begin with pixel-level motion detection following existing works (Zhang et al., 2024; Han et al., 2025). For each frame pair, we compute optical flow using off-the-shelf models (Wang et al., 2024a; Teed & Deng, 2020; An et al., 2025) ($\text{Flow}_{\text{optical}}$) and compare it against the flow induced by camera motion alone ($\text{Flow}_{\text{warp}}$). Regions where the L1 difference exceeds a specific threshold $\tau$ are marked as potentially dynamic:
$M_{i}^{\text{pixel}}=\mathds{1}\left[\|\text{Flow}_{\text{optical}}-\text{Flow}_{\text{warp}}\|_{1}>\tau\right],$ (6)
where $\mathds{1}[\cdot]$ denotes the element-wise indicator that returns 1 if the condition holds for a pixel and 0 otherwise, i.e., it flags pixels whose optical flow deviates from the warped flow by more than $\tau$.
However, pixel-level detection captures motion only at specific instants and misses complete object boundaries. We therefore propagate these sparse detections to full objects using SAM2 (Ravi et al., 2024), where we sample points from dynamic pixels in the first frame for prompts. Yet this approach still has limitations: static objects that begin moving in later frames may not be detected if they appear static initially.
Our solution employs backward tracking with CoTracker3 (Karaev et al., 2024) to aggregate motion evidence across the entire sequence. From the sampled points in each frame obtained from our pixel-level motion detection, we track these points from all frames back to $t=0$, capturing motions of objects that move at any point. These aggregated points are used to prompt the final SAM2 pass, producing complete object-level masks $M_{i}^{\text{obj}}$ that cleanly separate all dynamic content (Fig. 4-(b)). With the full dynamic mask, we can now obtain the static-only 3D geometry $\mathbf{P}_{\text{static}}$:
$\mathbf{P}_{\text{static}}=\bigcup_{i=1}^{L}\mathbf{P}_{i}\odot(1-M_{i}^{\text{obj}}).$ (7)
From the constructed static-only 3D geometry $\mathbf{P}_{\text{static}}$ with our proposed dynamic masking strategy, we now obtain the 3D scene memory:
$\mathcal{M}=(\hat{\mathbf{C}},\mathbf{P}_{\text{static}}),$ (8)
where we now explain how this 3D scene memory $\mathcal{M}$ can be used for scene-consistent camera-controllable video generation in the following section.
Figure 5: Dynamic masking. A three-stage pipeline refines dynamic region detection to produce complete object-level masks: (1) optical-flow differences detect pixel-level motion (Dynamic thresholding); (2) sample points from these regions for all frames and perform backward tracking (BW tracking) with CoTracker3 (Karaev et al., 2024) to aggregate motion evidence across all frames back to t=0t{=}0 (dynamic aggregation), capturing objects that move at any time; (3) propagate aggregated points in the first frame to the entire video using SAM2 (Ravi et al., 2024). The resulting dynamic masks cleanly separate moving elements (people, objects) from the static background, enabling construction of the static-only point cloud $\mathbf{P}_{\text{static}}$.
3.4 3D Scene Prompting
Having constructed the static-only 3D representation $\mathbf{P}_{\text{static}}$, rather than naïvely retrieving $T$ frames from the input video based on viewpoint similarity, we synthesize static-only spatial frames through the projection of $\mathbf{P}_{\text{static}}$. For each target camera pose $C_{t}\in\mathbf{C}$, we generate the corresponding spatial frame by projecting the static points from the most spatially relevant input frames:
$\text{Spatial}(t)=\Pi(K\cdot C_{t}\cdot\mathbf{P}_{\text{static}}^{(n)}),$ (9)
where $\mathbf{P}_{\text{static}}^{(n)}\subset\mathbf{P}_{\text{static}}$ contains points from the top-$n$ spatially adjacent input frames to $C_{t}$, $\Pi(\cdot)$ denotes perspective projection, and $K$ is the camera intrinsic matrix. The complete spatial conditioning becomes $\text{Spatial}(T)=\{\text{Spatial}(t)\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W\times 3}$, where spatial adjacency is calculated by field-of-view overlap.
This projection-based approach ensures only static content appears in conditioning while providing geometrically consistent views aligned to target poses. Notably, the static point cloud aggregates information from multiple viewpoints, potentially filling regions occluded by dynamic objects. These projected views serve as 3D scene prompts that provide explicit guidance about persistent scene structure, enabling precise camera control without additional encoding modules.
The projected views $\text{Spatial}(T)$ serve as what we term 3D scene prompts—they provide the model with explicit guidance about the persistent scene structure. By conditioning on both $\text{Temporal}(w)$ and $\text{Spatial}(T)$, our framework effectively enables scene-consistent camera-controllable video generation with computational efficiency while preserving the prior for high-quality video synthesis.
3.5 Training Dataset Construction
To train our model to effectively handle both static and dynamic scenes while leveraging spatial and temporal conditioning appropriately, we construct training data from diverse video sources with different preprocessing strategies based on scene characteristics.
Data processing.
Our training data comes from two primary sources: RealEstate10K (Zhou et al., 2018) containing static scenes, and OpenVid-1M (Nan et al., 2024), featuring dynamic content. For static videos from RealEstate10K, we directly extract 3D scene geometry without dynamic masking, as the scenes contain negligible motion. In contrast, for dynamic videos from OpenVid-1M, we employ a comprehensive dynamic masking pipeline explained in Section 3.3 to separate static and dynamic elements.
3D scene memory construction.
For effective training, we require videos of sufficient length to construct meaningful 3D scene context. To enable the model to handle arbitrary-length videos as input, we select long videos as the input video with $L>100$. As our video generation framework $\mathcal{F}(\cdot)$ generates $T$ frames, we designate the last $T$ frames as the ground truth future video that the model should generate, while frames before this window serve as input for constructing the 3D scene memory.
Using MegaSAM (Li et al., 2024), we process the entire video to estimate camera poses $\{\hat{\mathbf{C}}_{i}\}_{i=1}^{L}$ for all $L$ frames and reconstruct the 3D structure. We use frames $V_{\text{in}}\in\mathbb{R}^{(L-T)\times H\times W\times 3}$ (i.e., frames before the last $T$ frames) to construct the 3D scene memory, while the camera poses for the last $T$ frames $\{\hat{\mathbf{C}}_{i}\}_{i=L-(T-1)}^{L}$ are used during training to project the scene memory and generate spatial conditioning.
For static scenes, we directly use the reconstructed geometry from $V_{\text{in}}$. For dynamic scenes, we apply our dynamic masks to extract static-only geometry:
$\mathbf{P}_{\text{static}}=\bigcup_{i=1}^{L-T}\mathbf{P}_{i}\odot(1-M^{\text{obj}}_{i})$ (10)
The resulting 3D scene memory $\mathcal{M}=(\hat{\mathbf{C}},\mathbf{P}_{\text{static}})$ provides the spatial context for generating the future $T$ frames. During training, for each target frame $t$ in the last $T$ frames, we use its corresponding camera pose $\hat{\mathbf{C}}_{t}$ to project $\mathbf{P}_{\text{static}}$ from spatially adjacent viewpoints, creating the spatial conditioning Spatial(T)(T) that guides the generation process.
Spatial and temporal conditioning.
To train the model to utilize spatial conditioning effectively, it is crucial to incorporate both static and dynamic regions in our training data. The mixture of RealEstate10K and OpenVid-1M ensures the model tackles scenarios where spatial conditioning should dominate (static regions) and also maintains the prior of generating high-fidelity dynamic motions learned from large-scale pretraining.
For each training sample, we provide:
  • Temporal conditioning: The last $w=9$ frames from the input sequence, capturing recent motion dynamics.
  • Spatial conditioning: Projected views from the static 3D scene memory, synthesized by selecting the point clouds from the top-$n$ spatially adjacent viewpoints based on field-of-view overlap.
This dual conditioning strategy, trained on both static and dynamic content, enables the model to learn when to rely on spatial prompts for scene consistency versus temporal frames for motion continuity. The diversity in our training data—from completely static architectural scenes to videos with dynamic motion—ensures robust generalization to real-world scenarios where both static and dynamic elements coexist.
4 Experiments
4.1 Implementation Details
Model architecture.
We build upon CogVideoX-I2V-5B (Yang et al., 2024), extending its single-image conditioning to accept dual spatio-temporal inputs with minimal architectural changes. The key modification is repurposing the existing image conditioning channel to accept concatenated latents from both temporal frames and spatial projections. Specifically, we provide the last $w=9$ frames from $\mathbf{V}_{\text{in}}$ as temporal conditioning and $T$ projected views from the static point cloud as spatial conditioning. This enables the DiT backbone to remain entirely unchanged, preserving all pretrained video priors. Both conditions are encoded through the frozen 3D VAE and concatenated channel-wise such that $\mathbf{Z}_{\text{cond}}=\text{Concat}[\mathcal{E}(\text{Temporal}(w)),\mathcal{E}(\text{Spatial}(T))]$.
Fine-tuning.
We fully fine-tune the model for a total of 4K iterations with a batch size of 8 using 4 H100 GPUs, which required approximately 48 hours. We used the 16-bit Adam optimizer with a learning rate of $1\times 10^{-5}$, and adopted the same hyperparameter settings as those used in the training of CogVideoX (Yang et al., 2024). For the temporal sliding window, we provide the last 9 frames of the input video, setting $w=9$. For the projection of top-$n$ spatially adjacent views, we set $n=7$.
Experimental settings.
We evaluate our method across four key aspects: camera controllability, video quality, scene consistency, and geometric consistency. Since no prior work directly addresses scene-consistent camera-controllable video generation, we compare against two categories of baselines: (1) camera-controllable methods (CameraCtrl (He et al., 2024), MotionCtrl (Wang et al., 2024b), FloVD (Jin et al., 2025), AC3D (Bahmani et al., 2025)) for camera control and video quality metrics, and (2) DFoT (Song et al., 2025), which attempts scene-consistent camera-controllable generation, for spatial and geometric consistency metrics.
We primarily evaluate on 1,000 dynamic videos from DynPose-100K (Rockwell et al., 2025). For scene consistency evaluation, we additionally test on 1,000 static videos from RealEstate10K (Zhou et al., 2018), as static scenes provide clearer spatial consistency assessment.
Methods RealEstate10K DynPose-100K
[…]
DFoT Song et al. (2025) 18.3044 0.5960 0.3077 0.181164 12.1471 0.3040 0.4172 0.183202
3DScenePrompt (Ours) 20.8932 0.7171 0.2120 0.040843 13.0468 0.3666 0.3812 0.124189
Table 1: Evaluation of spatial and geometric consistency. We compare DFoT and Ours on the RealEstate10K (Zhou et al., 2018) and DynPose-100K (Rockwell et al., 2025) datasets. For spatial consistency, we evaluate PSNR, SSIM, and LPIPS on revisited camera trajectories, while for geometric consistency, we report the MEt3R (Asim et al., 2025) metric.
4.2 Scene-consistent video generation
Evaluation Protocol.
As mentioned in Section 3.1, one of the unique and key challenges in scene-consistent camera-controllable video generation is maintaining spatial consistency over extended durations. From a given input video, we evaluate spatial consistency by generating camera trajectories that revisit the viewpoints in the given video. By matching frames in the generated video and the input video that share the same viewpoint, we calculate PSNR, SSIM, and LPIPS. For RealEstate10K, we evaluate the whole image, whereas we only evaluate the static regions by masking out the dynamic regions for DynPose-100K. We also assess geometric consistency using Met3R (Asim et al., 2025), which measures multi-view alignment of generated frames under the recovered camera pose.
Results.
As shown in Tab. 1, 3DScenePrompt significantly outperforms DFoT across all metrics for both static and dynamic scenes. Most notably, our MEt3R evaluation error drops 77% (0.041 vs 0.181), demonstrating superior multi-view geometric alignment. While DFoT similarly tackles scene-consistent camera-controllable video generation through history guidance, their approach fails to maintain scene-consistency for long sequences due to memory constraints. In contrast, our dual spatio-temporal conditioning enables long-term scene-consistency without causing significant computational overhead. The qualitative comparisons shown in Fig. 6 also validate the effectiveness of our approach over DFoT.
Figure 6: Visualization of generated videos following trajectories that revisit early frames in the input video. We visualize and compare frames obtained from DFoT (Song et al., 2025) and Ours. We condition both DFoT and ours to generate a video that follows the camera trajectory shown in the top row (GT Future Video), which revisits a viewpoint in the input (Input Video). The comparison shows that ours shows much more consistent generation, whereas DFoT fails to generate scene-consistent frames mainly due to the limited number of frames it can condition on. Also, the comparison of the camera trajectory shows that our method more faithfully follows the given camera condition.
Figure 7: Visualization of scene-consistent camera-controllable video generation. Comparison of different methods  (Wang et al., 2024b; He et al., 2024; Jin et al., 2025; Bahmani et al., 2025) for generating videos from the same input (shown in Input Video) that follow the camera trajectory shown in the rightmost column (Camera). Our method best preserves scene consistency with the input video. Note the red-boxed regions: while the input video shows a white wall, competing methods either lose scene detail or fail to maintain the original scene structure. In contrast, our approach accurately remembers the white wall and maintains consistent scene elements throughout generation. In addition, when compared with the GT Future Video, ours best follows the camera condition, effectively verifying the strength of our framework for scene-consistent camera-controllable video generation.
4.3 Camera-Controllable Video Generation
Evaluation Protocol.
We employ the evaluation protocol of previous methods (He et al., 2024; Zheng et al., 2024; Jin et al., 2025) for the camera controllability. We provide an input image along with associated camera parameters for I2V models (He et al., 2024; Wang et al., 2024b; Jin et al., 2025) and solely provide camera parameters for the T2V model (Bahmani et al., 2025). For our model, we provide the last 9 frames of the input video together with the camera parameters. To evaluate how faithfully the generated video follows the camera condition, we estimate camera parameters from the synthesized video using MegaSAM (Li et al., 2024), and compare the estimated camera parameters against the condition camera trajectory $\mathbf{C}$.
As done in previous works (Hong et al., 2024b; a; Rockwell et al., 2025), the comparison between the estimated and input camera parameters is quantified using three metrics: mean rotation error (mRotErr), mean translation error (mTransErr), and mean error in the camera extrinsic matrices (mCamMC).
Table 2: Camera controllability evaluation.
Methods DynPose-100K
[…]
MotionCtrl Wang et al. (2024b) 3.5654 7.8231 9.7834
CameraCtrl He et al. (2024) 3.3273 9.5989 11.2122
FloVD Jin et al. (2025) 3.4811 11.0302 12.6202
AC3D Bahmani et al. (2025) 3.0675 9.7044 11.1634
DFoT Song et al. (2025) 2.3977 8.0866 9.2330
3DScenePrompt (Ours) 2.3772 7.4174 8.6352
For the generated video, we also assess video synthesis performance using the Fréchet Video Distance (FVD) (Skorokhodov et al., 2022) and seven metrics from VBench++ (Huang et al., 2024): subject consistency, background consistency, aesthetic quality, imaging quality, temporal flickering, motion smoothness, and dynamic degree.
Results.
We first evaluate camera controllability and compare our method with competitive baselines. As shown in Tab. 2, our approach consistently outperforms existing methods, indicating 3DScenePrompt is capable of generating videos with precise camera control. We then assess the overall video quality (Tab. 3) and provide qualitative comparisons (Fig. 7). As observed in Tab. 3, our method achieves the best generation quality across all metrics for dynamic video generation, which is further supported by the visual results in Fig. 7.
Table 3: Evaluation of video generation quality. We assess the quality of generated videos using FVD and VBench++ scores. For FVD, lower values indicate higher video quality. For VBench++ scores, higher values indicate better performance for all metrics. All VBench++ scores are normalized.
Methods DynPose-100K
FVD
Overall
Score
Subject
Consist
Bg
Consist
Aesthetic
Quality
Imaging
Quality
Temporal
Flicker
Motion
Smooth
Dynamic
Degree
MotionCtrl Wang et al. (2024b) 1017.4247 0.5625 0.5158 0.7093 0.3157 0.3149 0.8297 0.8432 0.7900
CameraCtrl He et al. (2024) 737.0506 0.6280 0.6775 0.8238 0.3736 0.3888 0.6837 0.6955 0.9900
FloVD Jin et al. (2025) 171.2697 0.7273 0.7964 0.8457 0.4722 0.5546 0.7842 0.8364 0.9900
AC3D Bahmani et al. (2025) 281.2140 0.7428 0.8360 0.8674 0.4766 0.5381 0.8020 0.8673 1.0000
3DScenePrompt (Ours) 127.4758 0.7747 0.8669 0.8727 0.4990 0.5964 0.8551 0.9260 1.0000
4.4 Ablation studies
Table 4: Ablation study on varying $n$.
Methods
Dynamic
mask $\mathcal{M}$
DynPose-100K
[…]
Ours ($n=1$) 13.0207 0.3732 0.3771 0.124773
Ours ($n=4$) 13.0382 0.3733 0.3758 0.124893
Ours ($n=L$) 13.0206 0.3631 0.3810 0.123507
Ours ($n=7$) 12.2304 0.3063 0.3821 0.134885
Ours ($n=7$) 13.0468 0.3666 0.3812 0.124189
We analyze two critical components of our framework: the dynamic masking strategy that separates static and dynamic elements, and the number of spatially adjacent frames $n$ retrieved for spatial conditioning. Tab. 4 demonstrates the impact of varying $n$ and the necessity of dynamic masking. Without dynamic masking (4th row), the model suffers significantly across all, showing a large drop of PSNR of approximately 0.8dB and also an increase of MEt3R error. This degradation occurs because unmasksed dynamic objects create ghosting artifacts when warped to new viewpoints, corrupting the spatial conditioning. Regarding the number of spatially adjacent frames, we find that performance stabilizes around $n=7$, with minimal improvements beyond this point, suggesting that 7 frames provide sufficient spatial context while maintaining computational efficiency.
5 Conclusion
In this work, we introduced 3DScenePrompt, a framework for scene-consistent camera-controllable video generation. By combining dual spatio-temporal conditioning with a static-only 3D scene memory constructed through dynamic SLAM and our dynamic masking strategy, we enable generating continuations from arbitrary-length videos while preserving scene geometry and allowing natural motion evolution. Extensive experiments demonstrate superior performance in camera controllability, scene consistency, and generation quality compared to existing methods. Our approach opens new possibilities for long-form video synthesis applications where maintaining both spatial consistency and precise camera control is essential.